Your golden retriever Maxwell barrels into the room and begins sliding on the freshly waxed floor. If he enters the room at a velocity of 6.0 m/s, and slides 6 meters in 2.2 seconds, what will his acceleration be?
1
Expert's answer
2020-12-21T10:38:20-0500

The distance "d = 6m" traveled under the constant acceleration "a" in time "t = 2.2s" is given with the formula:


"d = v_0t -\dfrac{at^2}{2}"

where "v_0 = 6m\/s" is the initial speed. Expressing the acceleration, obtain:


"a = \dfrac{2v_0}{t} - \dfrac{2d}{t^2}\\
a = \dfrac{2\cdot 6}{2.2} - \dfrac{2\cdot 6}{2.2^2} \approx 2.98m/s^2"

Answer. 2.98 m/s^2.